### v1

@@ 5.2 Equivalence between second-order stationary points
Subsequently, we turn to the second-order condition. Given an arbitrary $\eta\in\mathrm{T}_{\mathcal{M}}(X)$ and an associated $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$. By $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$, pick $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\eta=\mathrm{\bf L}_{Y}(v)$. According to the assumed inclusion (5.4), for any $\varepsilon>0$, there exist $v_{i}^{(j)}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\mathrm{\bf L}_{Y}(v_{i}^{(j)})\to\eta$, $u_{i}^{(j)}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v_{i}^{(j)})$, and
$s^{(j)}\in\lim_{i\to\infty}\big(\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+\operatorname{im}(\mathrm{\bf L}_{Y})\big)\ \ \text{such that}\ \ \|\,\zeta-\sum_{j=1}^{N}\mu_{j}s^{(j)}\,\|<\varepsilon,$ (5.5)
where the coefficients $\{\mu_{j}\}_{j=1}^{N}$ satisfy $\mu_{j}\geq 0$ and $\sum_{j=1}^{N}\mu_{j}=1$. For each $j$, choose a sequence $\{z_{i}^{(j)}\}_{i\in\mathbb{N}}\subseteq\operatorname{im}(\mathrm{\bf L}_{Y})$ realizing $s^{(j)}$ with $s^{(j)}=\lim_{i\to\infty}\big(\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+z_{i}^{(j)}\big)$. Since $Y$ is second-order stationary for (P-M), the condition (5.1) indicates that for any $(v_{i}^{(j)},u_{i}^{(j)})$, it holds that $\langle{\nabla\bar{f}(Y),u_{i}^{(j)}}\rangle+\langle{v_{i}^{(j)},\nabla^{2}\bar{f}(Y)[v_{i}^{(j)}]}\rangle\geq 0$. Applying the equality (5.3), and substituting $z^{(j)}_{i}\in\operatorname{im}(\mathrm{\bf L}_{Y})$ together with the first-order condition $\nabla f(X)\bot\operatorname{im}(\mathrm{\bf L}_{Y})$ yield that
$\langle{\nabla f(X),\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+z^{(j)}_{i}}\rangle+\left\langle\mathrm{\bf L}_{Y}(v_{i}^{(j)}),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v_{i}^{(j)})]\right\rangle\geq 0.$
Letting $i\to\infty$ in the inequality, we obtain $\langle{\nabla f(X),s^{(j)}}\rangle+\left\langle\eta,\nabla^{2}f(X)[\eta]\right\rangle\geq 0$. Taking the convex combination with the coefficients $\{\mu_{j}\}_{j=1}^{N}$ gives
$\Big\langle\nabla f(X),\,\sum_{j=1}^{N}\lambda_{j}s^{(j)}\Big\rangle+\langle\eta,\nabla^{2}f(X)[\eta]\rangle\geq 0.$
@@
We then turn to the “only if ” part. Firstly, suppose that $\operatorname{im}(\mathrm{\bf L}_{Y})\subsetneq\mathrm{T}_{\mathcal{M}}(X)$. Taking polars reverses the inclusion $(\mathrm{T}_{\mathcal{M}}(X))^{\circ}\subsetneq\operatorname{im}(\mathrm{\bf L}_{Y}))^{\circ}$, where we note that the strict inclusion still holds since $\operatorname{im}(\mathrm{\bf L}_{Y})$ is a linear space. Pick $w\in\mathrm{T}_{\mathcal{M}}(X)\setminus\operatorname{im}(\mathrm{\bf L}_{Y})$. Let $H=\mathcal{P}_{\operatorname{span}\{w\}}$ and $f(\tilde{X})=\tfrac{1}{2}\langle{\tilde{X}-X,H(\tilde{X}-X)}\rangle$ with $\mathcal{P}_{\operatorname{span}\{w\}}$ representing the projection matrices onto $\operatorname{span}\{w\}$. Then $\nabla f(X)=0$ and $\nabla^{2}f(X)=H$. By the computation (5.2), $\nabla\bar{f}(Y)=0$ and $\nabla^{2}\bar{f}(Y)=\mathrm{\bf L}_{Y}^{*}\circ H\circ\mathrm{\bf L}_{Y}\succeq 0$. Hence $Y$ is second-order stationary for (P-M). However, at $X$, $\langle\nabla f(X),w\rangle=0$ and $\langle{w,\nabla^{2}f(X)[w]}\rangle=-\|w\|^{2}<0$, violating the second-order condition for (P). Therefore, to guarantee the property $2\!\Rightarrow\!2$”, $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ must hold.
Denote by $\mathcal{C}$ the set on the right side of (5.4). Then, we suppose that the inclusion (5.4) fails: there exist a $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ (accordingly, $\eta:=\mathrm{\bf L}_{Y}(v)$) and a $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$ such that $\zeta\notin\mathcal{C}$. Since $\mathcal{C}$ is both closed and convex, we apply the separation theorem to obtain $w\in\mathcal{E}\setminus\{0\}$ such that
$\langle w,\zeta\rangle\ <\ 0\ \ \text{and}\ \ \langle w,s\rangle\ \geq\ 0\ \ \text{for all}\ s\in\mathcal{C}.$
Noticing that the subspace $\operatorname{im}(\mathrm{\bf L}_{Y})$ belongs to $\mathcal{C}$, we have $w\perp\operatorname{im}(\mathrm{\bf L}_{Y})$, and thus $\mathrm{\bf L}_{Y}^{*}(w)=0$. Let $H=\mathcal{P}_{\operatorname{im}(\mathrm{\bf L}_{Y})\cap\operatorname{span}\{\eta\}^{\perp}}$, and then, for all $v^{\prime}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, it holds that
$\langle w,\,\mathrm{\bf Q}_{Y,v^{\prime}}(u_{v^{\prime}})\rangle+\langle\mathrm{\bf L}_{Y}(v^{\prime}),H\circ\mathrm{\bf L}_{Y}(v^{\prime})\rangle\geq 0,\ \ \text{for all}\ u_{v^{\prime}}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v^{\prime}),$ (5.6)
since $\langle w,\cdot\rangle\geq 0$ on $\mathcal{C}$ and $\mathrm{\bf Q}_{Y,v^{\prime}}(u_{v^{\prime}})\in\mathcal{C}$ by definition of $\mathcal{C}$.
Let $f(\tilde{X})=\langle w,\tilde{X}-X\rangle+\tfrac{1}{2}\langle{\tilde{X}-X,H(\tilde{X}-X)}\rangle$, which reveals that $\nabla f(X)=w$ and $\nabla^{2}f(X)=H$. By the computations (5.2), (5.3), and (5.6), we have
$\nabla\bar{f}(Y)=\mathrm{\bf L}_{Y}^{*}(w)=0,\ \ \text{and}\ \ \langle{\nabla\bar{f}(Y),u_{v^{\prime}}}\rangle+\langle{v^{\prime},\nabla^{2}\bar{f}(Y)[v^{\prime}]}\rangle\geq\ 0$
for all $v^{\prime}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ and $u_{v^{\prime}}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v^{\prime})$. Therefore, $Y$ is second-order stationary for (P-M). On the other hand, at $X$, we have $\langle\nabla f(X),\eta\rangle=\langle w,\mathrm{\bf L}_{Y}(v)\rangle=\langle{\mathrm{\bf L}_{Y}^{*}(w),\eta}\rangle=0$, and for the associated $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$,
@@
$\langle\nabla f(X),\zeta\rangle+\langle{\eta,\nabla^{2}f(X)[\eta]}\rangle=\underbrace{\langle w,\zeta\rangle}_{<\,0}+\underbrace{\langle{\eta,H(\eta)}\rangle}_{=0}\ <\ 0,$
@@
which violates the second-order optimality condition on $\mathcal{M}$. This contradicts the property $2\!\Rightarrow\!2$”, thereby showing the necessity of the inclusion (5.4).

### v2

@@ 5.2 Equivalence between second-order stationary points
Subsequently, we turn to the second-order condition. Given an arbitrary $\eta\in\mathrm{T}_{\mathcal{M}}(X)$ and an associated $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$. By $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$, pick $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\eta=\mathrm{\bf L}_{Y}(v)$. According to the assumed inclusion (5.6), for any $\varepsilon>0$, there exist $\{(s^{(j)},b^{(j)},\tau^{(j)})\}_{j=1}^{N}\subseteq\mathcal{Q}_{Y}(v)\times\mathcal{Q}_{Y}(0)\times\varGamma_{Y}$, together with the coefficients $\{\mu_{j}\}_{j=1}^{N}$ satisfying $\mu_{j}\geq 0$ and $\sum_{j=1}^{N}\mu_{j}=1$, such that
$\|\,\zeta-\sum_{j=1}^{N}\mu_{j}(s^{(j)}+b^{(j)}+\tau^{(j)})\,\|<\varepsilon.$ (5.7)
For all $j=1,2,\ldots,N$, applying Lemmas 8 and 9, we have
$\langle{\nabla f(X),s^{(j)}}\rangle+\left\langle\eta,\nabla^{2}f(X)[\eta]\right\rangle\geq 0,\ \ \langle\nabla f(X),b^{(j)}\rangle{\geq 0},\ \ \text{and}\ \langle{\nabla f(X),\tau^{(j)}}\rangle=0.$
Taking the convex combination with the coefficients $\{\mu_{j}\}_{j=1}^{N}$ gives
$\Big\langle\nabla f(X),\,\sum_{j=1}^{N}\mu_{j}(s^{(j)}+b^{(j)}+\tau^{(j)})\Big\rangle+\langle\eta,\nabla^{2}f(X)[\eta]\rangle\geq 0.$
@@
(Necessity) We then turn to the “only if ” part. Firstly, suppose that $\operatorname{im}(\mathrm{\bf L}_{Y})\subsetneq\mathrm{T}_{\mathcal{M}}(X)$. Taking polars reverses the inclusion $(\mathrm{T}_{\mathcal{M}}(X))^{\circ}\subsetneq\operatorname{im}(\mathrm{\bf L}_{Y}))^{\circ}$, where we note that the strict inclusion still holds since $\operatorname{im}(\mathrm{\bf L}_{Y})$ is a linear space. Pick $w\in\mathrm{T}_{\mathcal{M}}(X)\setminus\operatorname{im}(\mathrm{\bf L}_{Y})$. Let $w_{\bot}:=w-\mathcal{P}_{\operatorname{im}(\mathrm{\bf L}_{Y})}(w)$, and the operator $H:z\mapsto-\langle{w_{\bot},z}\rangle w_{\bot}$. Define $f(\tilde{X})=\tfrac{1}{2}\langle{\tilde{X}-X,H(\tilde{X}-X)}\rangle$. Then $\nabla f(X)=0$ and $\nabla^{2}f(X)=H$. By the computation (5.3), $\nabla\bar{f}(Y)=0$ and $\langle{v,\nabla^{2}\bar{f}(Y)[v]}\rangle=\langle\mathrm{\bf L}_{Y}(v),H\circ\mathrm{\bf L}_{Y}(v)\rangle=0$ for all $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Hence $Y$ is second-order stationary for (P-M). However, at $X$, $\nabla f(X)=0$ and $\langle{w,\nabla^{2}f(X)[w]}\rangle=-\|w_{\bot}\|^{4}<0$, violating the second-order condition for (P). Therefore, to guarantee the property $2\!\Rightarrow\!2$”, $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ must hold.
Then, we assume that condition (5.6) fails for some $v$. Let $\eta=\mathrm{\bf L}_{Y}(v)$. If $\eta=0$, we have
$\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{M}}(X;0)=\mathrm{T}_{\mathcal{M}}(X)=\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathcal{Q}_{Y}(0),$
where the second equality holds from (giorgi2010overviewsectengent, Proposition 9), and the last inclusion comes from Lemma 7. Therefore, the inclusion (5.6) always holds for $v\in\operatorname{ker}(\mathrm{\bf L}_{Y})$, and thus we assume that (5.6) fails for some $v$ with $\|\eta\|>0$ in the following discussion. Then there exists $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$ outside the closed convex set $\mathcal{C}_{v}:=\overline{\mathrm{conv}}(\mathcal{Q}_{Y}(v)+\mathcal{Q}_{Y}(0)+\varGamma_{Y})$. By the strict separation theorem boyd2004convex, there exists a vector $w\in\mathcal{E}$, a threshold $\kappa\in\mathbb{R}$, and a gap $\delta>0$ such that:
$\langle{w,\zeta}\rangle\leq\kappa-\delta<\kappa:=\inf_{p\in\mathcal{C}_{v}}\langle{w,p}\rangle.$ (5.8)
Since $\mathcal{Q}_{Y}(0)$ is a cone (see Lemma 7), we have $\langle{w,u}\rangle\geq 0$ for all $u\in\mathcal{Q}_{Y}(0)$. Noticing that the subspace $\operatorname{im}(\mathrm{\bf L}_{Y})$ belongs to $\mathcal{Q}_{Y}(0)$, we have $w\perp\operatorname{im}(\mathrm{\bf L}_{Y})$. Additionally, since $c\varGamma_{Y}\subseteq\varGamma_{Y}$ for any scaling $c\in\mathbb{R}$, we have $w\perp\varGamma_{Y}$. In summary, it holds that
$\langle{w,u_{0}}\rangle\geq 0,\ \ \langle{w,u_{1}}\rangle=0,\ \ \text{and}\ \langle{w,u_{2}}\rangle=0,$ (5.9)
for all $u_{0}\in\mathcal{Q}_{Y}(0)$, $u_{1}\in\operatorname{im}(\mathrm{\bf L}_{Y})$, and $u_{2}\in\varGamma_{Y}$.
We then construct the following function,
$f_{\rho}(\tilde{X}):=\langle{w,\tilde{X}-X}\rangle+\frac{1}{2}\langle{\tilde{X}-X,(H_{\mathrm{shift}}+\rho H)(\tilde{X}-X)}\rangle,$
where $H=I-\mathcal{P}_{\operatorname{span}\{\eta\}}$ and $H_{\mathrm{shift}}=\frac{-\kappa+\delta/2}{\|\eta\|^{2}}\mathcal{P}_{\operatorname{span}\{\eta\}}$. The Euclidean derivatives of $f_{\rho}$ at $X$ are provided below,
$\nabla f_{\rho}(X)=w,\ \ \ \text{and}\ \ \ \nabla^{2}f_{\rho}(X)=H_{\mathrm{shift}}+\rho H.$
First, we observe that $X$ is not a second-order stationary point of $f_{\rho}$. Specifically, along the direction $(\eta,\zeta)$, the second-order condition fails:
@@
$\langle{\nabla f_{\rho}(X),\zeta}\rangle+\langle{\eta,\nabla^{2}f_{\rho}(X)[\eta]}\rangle\!=\!\langle{w,\zeta}\rangle+\langle{\eta,H_{\mathrm{shift}}(\eta)}\rangle\!\leq\!(\kappa-\delta)+(-\kappa+\delta/2)=-\delta/2\!<\!0.$
Second, we aim to show that there exists a $\rho>0$ such that $Y\in\overline{\mathcal{M}}$ is a second-order stationary point for the objective $\bar{f}_{\rho}=f_{\rho}\circ\phi$. Given any direction $d\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, we consider the following function induced by the Hessian of $\bar{f}_{\rho}$ at $Y$:
$\mathcal{H}_{\rho}(d):=$ $\ \langle{\nabla\bar{f}_{\rho}(Y),u_{d}}\rangle+\langle{d,\nabla^{2}\bar{f}_{\rho}(Y)[d]}\rangle$
$=$ $\ \langle{\nabla{f}_{\rho}(X),\mathrm{\bf Q}_{Y,d}(u_{d})}\rangle+\langle{\mathrm{\bf L}_{Y}(d),\nabla^{2}{f}_{\rho}(X)[\mathrm{\bf L}_{Y}(d)]}\rangle$ (5.10)
$=$ $\ \underbrace{\langle{w,\mathrm{\bf Q}_{Y,d}(u_{d})}\rangle+\langle{\mathrm{\bf L}_{Y}(d),H_{\mathrm{shift}}\circ\mathrm{\bf L}_{Y}(d)}\rangle}_{\text{A quadratic form}\ E(d,d)}+\rho\|H(\mathrm{\bf L}_{Y}(d))\|^{2},$ (5.11)
where (5.10) comes from (5.4). Since $w\perp\operatorname{im}(\mathrm{\bf L}_{Y})$, the identity (5.2) implies that the term $\langle{w,\mathrm{\bf Q}_{Y,d}(u_{d})}\rangle=\langle{w,\mathrm{D}\phi_{Y}(\mathrm{II}_{Y}(d,d))+\mathrm{D}^{2}\phi_{Y}[d,d]}\rangle$. Therefore, the value of $\langle{w,\mathrm{\bf Q}_{Y,d}(u_{d})}\rangle$ is independent of the choice of representative $u_{d}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;d)$, and the function $\mathcal{H}_{\rho}(d)$ and the quadratic form $E(d,d)$ given in (5.11) are well-defined.
Let $K:=\ker(H\circ\mathrm{\bf L}_{Y})=\mathrm{\bf L}_{Y}^{-1}(\operatorname{span}(\eta))$. We present two properties of $E(d,d)$:
  1. 1)
    Given $d\in\ker(\mathrm{\bf L}_{Y})$, it holds that $\mathrm{\bf Q}_{Y,d}(u_{d})\subseteq\mathcal{Q}_{Y}(0)$. By $\mathrm{\bf L}_{Y}(d)=0$ and the separation property (5.9), we have $E(d,d)=\langle{w,\mathrm{\bf Q}_{Y,d}(u_{d})}\rangle\geq 0$.
  2. 2)
[…]
In summary, $E(d,d)\geq 0$ for all $d\in K$. Let $K_{0}=\{d\in K\mid E(d,d)=0\}$, and the property 2) indicates that $K_{0}\subseteq\ker(\mathrm{\bf L}_{Y})$. More importantly, we find that
$E(d_{0},d)=0,\ \ \ \text{for all}\ d_{0}\in K_{0},\ \text{and}\ d\in\mathrm{T}_{\overline{\mathcal{M}}}(Y),$ (5.12)
the derivation of which is deferred to the end of this proof. Now, let $K_{+}$ be the orthogonal complement of $K_{0}$ in $K$, and let $K^{\bot}$ be the orthogonal complement of $K$ in $\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, that is, $\mathrm{T}_{\overline{\mathcal{M}}}(Y)=K_{0}\oplus K_{+}\oplus K^{\bot}$. Then, there exists a $\lambda>0$ such that $E(d_{+},d_{+})\geq\lambda\|d_{+}\|^{2}$ for all $d_{+}\in K_{+}$.
We claim that there exists a $\rho>0$ such that $\mathcal{H}_{\rho}(d)\geq 0$ for all $d\in\mathcal{E}$; otherwise, we can find a sequence $(\rho_{k})_{k\in\mathbb{N}}$ with $\rho_{k}\to\infty$ and unit vectors $d_{k}$ such that $\mathcal{H}_{\rho_{k}}(d_{k})<0$ for every $k$. Decompose $d_{k}=d_{k,0}+d_{k,+}+\ell_{k}$, where $d_{k,0}\in K_{0}$, $d_{k,+}\in K_{+}$, and $\ell_{k}\in K^{\perp}$. The penalty term satisfies $\|H(\mathrm{\bf L}_{Y}(d_{k}))\|=\|H(\mathrm{\bf L}_{Y}(\ell_{k}))\|\geq\sigma\|\ell_{k}\|$ for some $\sigma>0$. Using $E(d_{k,0},\cdot)=0$, the condition $\mathcal{H}_{\rho_{k}}(d_{k})<0$ becomes $E(d_{k,+}+\ell_{k},d_{k,+}+\ell_{k})+\rho_{k}\sigma^{2}\|\ell_{k}\|^{2}<0$. We then expand and bound the terms to obtain
$\lambda\|d_{k,+}\|^{2}-2\|E\|\|d_{k,+}\|\|\ell_{k}\|+(\rho_{k}\sigma^{2}-\|E\|)\|\ell_{k}\|^{2}<0.$
If $\ell_{k}=0$, then $\lambda\|d_{k,+}\|^{2}<0$ leads to a contradiction. If $\ell_{k}\neq 0$, dividing by $\|\ell_{k}\|^{2}$ yields the following quadratic inequality with $t_{k}=\|d_{k,+}\|/\|\ell_{k}\|$:
$\lambda t_{k}^{2}-2\|E\|t_{k}+(\rho_{k}\sigma^{2}-\|E\|)<0.$ (5.13)
The minimum value of the left side is $\rho_{k}\sigma^{2}-\|E\|-\frac{\|E\|^{2}}{\lambda}$. As $\rho_{k}\to\infty$, this minimum becomes positive, making the inequality (5.13) impossible. Thus, there indeed exists a sufficiently large $\rho$ such that $\mathcal{H}_{\rho}(d)\geq 0$ for all $d\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Consequently, we have
$\langle{\nabla\bar{f}_{\rho}(Y),u_{d}}\rangle+\langle{d,\nabla^{2}\bar{f}_{\rho}(Y)(d)}\rangle={\mathcal{H}_{\rho}(d)}\geq 0,$
@@
for all $d\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ and $u_{d}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;d)$. Therefore, $Y$ is second-order stationary for (P-M), while $X$ is not second-order stationary for (P). This contradicts the property $2\!\Rightarrow\!2$”, thereby showing the necessity of the inclusion (5.6).
Derivation of (5.12) In fact, taking into account the definition of $E(d,d)$ (5.11), the property $w\perp\operatorname{im}(\mathrm{\bf L}_{Y})$ (5.9), and the equality (5.2), we have
$E(d,d)=\langle{w,\mathrm{D}\phi_{Y}(\mathrm{II}(d,d))+\mathrm{D}^{2}\phi_{Y}[d,d]}\rangle+\langle{\mathrm{\bf L}_{Y}(d),H_{\mathrm{shift}}\circ\mathrm{\bf L}_{Y}(d)}\rangle.$
Then, $\mathrm{\bf L}_{Y}(d_{0})=0$ reveals that
$E(d_{0},d)$ $=\langle{w,\mathrm{D}\phi_{Y}(\mathrm{II}(d_{0},d))+\mathrm{D}^{2}\phi_{Y}[d_{0},d]}\rangle$
$=\langle{w,\mathcal{P}_{\hat{\mathrm{N}}_{\mathcal{M}}(X)}\left({\mathrm{D}\phi_{Y}(\mathrm{II}(d_{0},d))+\mathrm{D}^{2}\phi_{Y}[d_{0},d]}\right)}\rangle,$
where the last equality holds from the separation properties $w\in(\operatorname{im}(\mathrm{\bf L}_{Y}))^{\circ}=(\mathrm{T}_{\mathcal{M}}(X))^{\circ}=\hat{\mathrm{N}}_{\mathcal{M}}(X)$. Finally, we employ the definition of the set $\varGamma_{Y}$ (5.5) and the property $w\perp\varGamma_{Y}$ (5.9) to conclude $E(d_{0},d)=0$ for all $d_{0}\in K_{0}$ and $d\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$.